Define binding energy. Calculate the binding energy per nucleon of Fe. 56
26 Given M(56 Fe) = ( Fe) 55.934932 u,mn =1.008665 u and mp=1.007825 u.
26
1
Expert's answer
2019-02-19T09:52:07-0500

The total mass of free nucleons equals to mn = 26*1.007825u + 30*1.008665u = u(26.20345 + 30.25995) = 56.4634u; respectively the mass defect is mD = u(56.4634 - 55.934932) = 0.528468u. The binding energy E = mDc2 = 0.528468*1.660540*10-27*8.987552*1016 = 7.886957 *10-11 (J) or 492.2 MeV. The binding energy per nucleon is En = E/56 = 8.79 (MeV/n).


The binding energy E = 7.886957 *10-11 J.

The binding energy per nucleon En = 8.79 MeV/n.